OpenStudy (anonymous):

\[\int\limits_{2}^{\infty}te^{-st}dt\]where "s" is some constant

OpenStudy (anonymous):

Looks like Laplace transform.

zepdrix (zepdrix):

Hmm is this coming from like.. part of a Laplace Transform of a piecewise?

OpenStudy (anonymous):

Yes sir it is

zepdrix (zepdrix):

oh cool c:

OpenStudy (anonymous):

Use integration by parts.

OpenStudy (anonymous):

didn't u get it last time ?

OpenStudy (anonymous):

I'll post the original question

OpenStudy (anonymous):

\(\large u=t\), \( \large dv = e^{-st}dt\)

OpenStudy (anonymous):

\(\\ \huge \color{red}{\int e^x[f(x)+f’(x)]dx=e^xf(x)+c} \\\)

OpenStudy (anonymous):

here f(t) =t+1

OpenStudy (anonymous):

Using the definition to determine the Laplace transform if the given function f(t)= {0, 0<t<2 {t, 2<t

zepdrix (zepdrix):

So you were able to figure out the first piece alright? Just having trouble with the other piece since it's by parts I guess ?:D

OpenStudy (anonymous):

@hartnn told me that @gohangoku58 I'm having trouble seeing how it applies

OpenStudy (anonymous):

\(\\ \huge \color{red}{\int e^{-st}[(t+1)-1]dx=e^{-st}(t+1)/(-s)+c} \\\)

OpenStudy (anonymous):

ya, its been a couple years since I had to use parts. I'd like to see it done both methods shown above.

OpenStudy (anonymous):

*dt

OpenStudy (anonymous):

@gohangoku58 ahh, thanks. I forgot about subtracting 1. But if f(t) = t + 1 f'(t) = 1

OpenStudy (anonymous):

ohh... then it will be t-1

OpenStudy (anonymous):

f(t) = t-1

OpenStudy (anonymous):

sorry :\

zepdrix (zepdrix):

 This would be the set up by parts, just in case you wanted to see it :O hopefully I didn't skip too many steps there.